OpenStudy (anonymous):

there are 155 people at the basketball game. tickets for the game are 2.50 for students and 4 for adults.if the total money received for the admission was 492.50, how many studenta and adults attended the game

OpenStudy (anonymous):

it dosent say it

OpenStudy (anonymous):

let there be s students and a adults at the game

OpenStudy (anonymous):

then we have a total of 155 people at the game so a + s = 155

OpenStudy (anonymous):

and also (2.5s) gives tickets for students and (4a) gives tickets for adults so total money becomes 2.5s + 4a = 492.5

OpenStudy (anonymous):

now solve a + s = 155 and 2.5s + 4a = 192.5 u'll get the answer

OpenStudy (anonymous):

thank u so much for ur help

OpenStudy (anonymous):

happy to help ;)

OpenStudy (anonymous):

(: